# Quantum heat engine in the optomechanical system with mechanical parametric drive

### v1

@@ I Introduction
Quantum thermodynamics [9] is an extension of the standard thermodynamics that the physical objects with the inclusion of quantum effects and usually in non-equilibrium situations. This rapidly growing research area not only provides the fundamental viewpoint on the understanding of quantum systems but also has great potential for nano-scale applications. As one of the important topics, quantum heat engines (QHEs) are important devices that convert heat to mechanical work for quantum systems [33, 6, 15, 34, 35]. Because of the quantum effects of the working substance or the heat baths, QHEs have unusual properties. For the cyclic heat engines to generate mechanical output work in a thermodynamical cycle, the efficiency of QHEs will surpass the standard limit defined in the classical thermodynamics [10, 11, 15]. While for the steady-state heat engines that transfer energies from the hot baths to the cold baths, the QHE suggests some effects violated the classical laws of thermodynamics [13, 14], although it is correct in the quantum mechanical viewpoint.
@@
To design a quantum heat engine with non-zero power output, the finite-time thermodynamical processes will be taken into consideration for the optimization of the QHE. The finite-time QHE has been discussed for the finite-time adiabatic processes [24, 25], and by the technique of shortcut-to-adiabatically (STA) [5, 22, 23]. These two strategies lead to higher efficiencies and non-zero power.
In this paper, we have proposed a quantum harmonic heat engine constructed in an optomechanical system with an external mechanical parametric drive [3]. Unlike the other studies for QHEs in the optomechanical systems [27, 26, 28, 29], the cavity is chosen as working substance. In particular, the mechanical mode has been considered as the quantum object that provides external energies, which is equivalent to the fuel as the component for classical machines. The heat flux for the total system in the cavity thermalization process is discussed, and the cyclic work extraction can be achieved from a single physical thermal reservoir coupled to the cavity. Due to the interaction of optomechanics, non-equilibrium behavior is induced, resulting in this counter-intuitive phenomena [12]. We then construct the Otto-type cycle for the heat engine by modulating the parametric driving strength [4]. Finally, the extra heat absorption induced by the optomechanical interaction in the work generation strokes has been considered, and the corresponding thermal efficiencies within and without the STA method, are analyzed.
@@
[…]
@@
where $H_{a}=\Delta(a^{\dagger}a+1/2)$, $H_{b}=\omega_{M}(b^{\dagger}b+1/2)$, $\Delta$ is the cavity detuning, $\omega_{M}$ is the intrinsic frequency for mechanical mode, $G$ is the collective coupling, $\lambda$ indicates the parametric drive strength. The cavity mode is chosen as the quantum harmonic heat engine, with the detuning changing as work producing. Here the cavity detuning $\Delta=\omega_{c}-\omega_{L}$ can be tuned in situ. The dynamics of the total optomechanical system can be described by the following master equation:
@@
$\begin{split}\mathcal{L}_{\text{diss}}\rho&=\kappa(\bar{n}_{c}+1)D[a]\rho+\kappa\bar{n}_{c}D[a^{\dagger}]\rho\\ &+\gamma(\bar{n}+1)D[b]\rho+\gamma\bar{n}D[b^{\dagger}]\rho.\end{split}$ (3)
Based on the second law of thermodynamics, a heat engine must be in contact with at least two thermal baths, with different thermal occupations $\bar{n}_{c}$ and decay rate $\kappa$ in corresponding thermodynamical processes. However, if an auxiliary term that interacts with the heat engine is considered, the interaction provides an additional excitation for the heat engine, which is equivalent to an effective hot (cold) reservoir only coupled with the heat engine. In our work, the optomechanical interaction and the mechanical parametric drive would excite the cavity into higher occupations at the non-equilibrium steady-states, and thus produce an effective hot thermal bath for the heat engine. Therefore, the mechanical mode with parametric drive here fuels the cavity, and the interaction term transports the energy from the fuel to the engine.
II.1 Heat flux within the parametric drive
As mentioned above, the mechanical mode acts as quantum fuel which provides effective hot reservoirs. Meanwhile, the cavity extracts heat from the mechanical mode via the optomechanical interaction. Naturally, the utilization efficiency of the energy of the quantum fuel, which can be defined as the ratio of heat flux transforming into the heat engine to the total heat flux injected into the mechanical mode, is one of the characteristics of quantum heat engines.
Based on the energy conservation, the heat flux for both the cavity and the mechanical mode has the form [2]
@@ II.1 Heat flux within the parametric drive
The detailed derivation is given in Appendix A, here $\dot{Q}_{x}$ indicates the different contribution for the heat flux. As illustrated in Fig. 1, $\dot{Q}_{\rm{PD}}$ is the energy injection via mechanical parametric drive, $\dot{Q}_{b\rightarrow a}$ indicates heat flux transferred from the fuel to the heat engine, $\dot{Q}^{\uparrow}_{i}(i=a,b)$ is the heating flux from the reservoir, $\dot{Q}^{\downarrow}_{i}{(i=a,b)}$ is the dissipating flux. The exact form of these heat flux terms can be derived from the master equation(Eq. (2)),
@@
The definition of utilization efficiency of energy for the mechanical mode is
@@
and the steady state occupation of photon mode is defined as $\bar{n}_{h}=\langle a^{\dagger}a\rangle_{ss}$. The steady-state utilization efficiency of energy produced from the quantum fuel, and the effective thermal occupations for the cavity, are plotted in Fig. 1. Compared with the standard optomechanical interaction case ($\lambda=0$), the external mechanical driven term generates additional energy injection, which supplies more energy than just extracted from the mechanical thermal bath. From the view of quantum thermodynamics, the larger steady-state occupation for the cavity indicates higher inner energy differences, which reflects more heat absorption or work production. On the other hand, following the discussion of heat flux for the total system, the larger parametric drive strength leads to the higher utilization efficiency of energy, thus more heat is transmitted from the mechanical mode into the cavity. Therefore, the cavity has a stronger ability to extract heat from the additional effective heat bath. Both of the above qualities indicate that in this optomechanical model, the cavity is a suitable candidate for heat engines used to produce mechanical output work, while the mechanical mode, which injects the additional heat into the engine and creates an effective hot thermal reservoir, has been acting as the quantum fuel.
@@ II.2 Quantum Otto-type Cycle
A typical Otto cycle is constructed by two isentropic and two adiabatic processes. In the isentropic process, the engine is in contact with the hot(cold) thermal bath and has the energy exchange, while in the adiabatic process, the engine does not have thermal contact with the environment and just produces work. However, in quantum scenarios, the adiabatic processes not only satisfy the thermal adiabaticity that the system has zero thermal exchange but also follow the quantum adiabatic dynamics [6, 10, 34]. Thus, we call the former case quantum Otto-type cycle.
@@
Figure 2: (a) The quantum Otto cycle for the cavity by tuning the detuning to produce the extra work output, and turning the parametric drive on and off to "heating" and "cooling" the cavity. (b) Amplitude and detuning for the laser drive in an Otto-type cycle.
(1). Isentropic compression($\rho_{1}\rightarrow\rho_{2}$), the cavity detuning $\Delta(t)$ and the amplitude of laser drive $A$ will be changed from ($\Delta_{l}$, $A_{l}$) to ($\Delta_{h}$, $A_{h}$) simultaneously during time $\tau$ to fix the classical part of the cavity mode. During this stroke, $\lambda=0$.
@@
To design the two isentropic strokes, the evolution time $\tau\ll 1/\kappa$ is achieved to neglect the energy dissipation. In the weak interacting $G\ll\omega_{M}$ and resolved sideband $\kappa\ll\omega_{M}$ regimes, the optomechanical interaction would only be resonantly enhanced. Hence, in the large detuned case, $|\Delta_{l}-\Delta_{h}|\gg\kappa$, the steady state for cavity in the cold isochoric stroke would almost be the thermal state with average occupation $\bar{n}_{c}$, and the optomechanical interaction term can be regarded as perturbations in analyzing the two isentropic strokes.
We first neglect the optomechanical interaction in the two isentropic strokes. The work production and heat absorption in a quantum Otto cycle can be defined with the two-point measurement approach, in which the energy change for the heat engine is obtained by projective measurement at the beginning and the end of each strokes [32]. The joint probability distribution for an Otto cycle has the form [20]
@@
Here $P^{\tau}_{n\rightarrow m}$ and $P^{\tau}_{k\rightarrow j}$ indicate the transition probabilities from initial eigenstates $n(k)$ to the final eigenstates $m(j)$ for the isentropic strokes, and $P_{n}(N)=(1/(1+N))(N/(1+N))^{n}$ is the probability distribution for the thermal states with occupation $N$.
For the calculation of average net work output and heat absorbed per cycle, the characteristic function of the joint probability distribution is
[…]
@@
$\langle W\rangle_{Q^{*}}$ $=-i\frac{\partial\rm{ln}\chi(\emph{u},\emph{v})}{\partial u}\mid_{u=v=0},$ (12a)
$\langle Q\rangle_{Q^{*}}$ $=-i\frac{\partial\rm{ln}\chi(\emph{u},\emph{v})}{\partial v}\mid_{u=v=0}.$ (12b)
@@
$\eta_{\rm{th}}=1-\frac{\Delta_{l}}{\Delta_{h}}\frac{Q^{*}(\bar{n}_{h}+1/2)-(\bar{n}_{c}+1/2)}{(\bar{n}_{h}+1/2)-Q^{*}(\bar{n}_{c}+1/2)}.$ (13)
@@ III Work generation and efficiencies
The quantum Otto engine reaches its maximal thermal efficiency when the isentropic strokes evolve quantum adiabaticity($Q^{*}=1$), with which the adiabatic Otto efficiency $\eta_{\rm{Qtto}}=1-\Delta_{l}/\Delta_{h}$. However, the quantum adiabatic condition requires the dynamical evolution of the system to be slow enough that the non-adiabatic transitions between different eigenstates can be neglected, and the system follows the instantaneous eigenstates in time. Therefore, even if the quantum adiabatic processes would have maximal efficiency, the power for the total work generation is zero. To obtain the efficiency close to the adiabatic case in finite time, the shortcut-to-adiabaticity(STA) technique [23, 5] has been considered.
@@ III.1 Extra heat produced by optomechanical interaction
Figure 4: Efficiencies for different controlling parameters under fixed $\tau$. (a). Efficiencies for different detunings with a fixed ratio of $\Delta_{l}/\Delta_{h}=1/2$, where the mechanical parametric driving strength $\lambda=0.4\kappa$; (b). Efficiencies vs $\lambda$ with the same ratio of $\Delta_{l}/\Delta_{h}=1/2$ and fixed $\Delta_{h}=20\kappa$. In both plots, the evolution time has been chosen as $\kappa\tau=0.1$. Other parameters for both plots are chosen as $\bar{n}_{c}=0.01$, $\bar{n}=100$, $G=\kappa$, $\gamma=0.01\kappa$ and $\omega_{M}=\Delta_{h}$.
The thermal efficiency within the optomechanical interaction $\eta^{\prime}_{\text{th}}$ can be easily obtained by the above procedure. The efficiencies for the fixed evolution time $\tau$ are plotted in Fig. 4. In both plots, the nonadiabatic efficiencies are calculated by Eq. (13), the modification for the heat absorption induced by the optomechanical interaction is added to the total heat absorption, and the detuning was tuned following the Eq. (16). The efficiencies for the STA methods that neglect and include the extra heat terms are calculated by Eq.(18) and Eq. (22), respectively. It has been noted that as $\Delta_{h}$ and $\Delta_{l}$ increase, the STA Hamiltonian(Eq. (14)) has less contribution to the system evolution, the system evolves near the instantaneous eigenstates once the boundary conditions are fulfilled. Meanwhile, because the contribution to the total heat absorption in a thermodynamic cycle has three parts, heat absorbing from the effective hot thermal reservoir $\langle Q\rangle$, the counter-diabatic driving $\langle H^{c/e}_{\rm{STA}}\rangle_{\tau}$ and the extra heat from the optomechanical interaction(Eq. (21)), the contribution for the latter two terms is mainly determined by the evolution time for the two strokes, thus the influence of these two parts decreases as mechanical parametric driving strength increases.

### v2

@@ I Introduction
Quantum thermodynamics [1] is an extension of standard thermodynamics that the physical objects with the inclusion of quantum effects and usually in non-equilibrium situations. This rapidly growing research area not only provides the fundamental viewpoint on the understanding of quantum systems but also has great potential for nano-scale applications. As one of the important topics, quantum heat engines (QHEs) are important devices that convert heat to mechanical work for quantum systems [2, 3, 4, 5, 6]. Because of the quantum effects of the working substance or the heat baths, QHEs have unusual properties. For the cyclic heat engines to generate mechanical output work in a thermodynamic cycle, the efficiency of QHEs will surpass the standard limit defined in classical thermodynamics [7, 8, 4]. While for the steady-state heat engines that transfer energies from the hot baths to the cold baths, the QHE suggests that certain effects may appear to violate classical thermodynamic laws [9, 10]. However, these effects are consistent with the principles of quantum mechanics.
@@
To design a quantum heat engine with non-zero power output, the finite-time thermodynamical processes will be taken into consideration for the optimization of the QHEs. The finite-time QHEs has been discussed for the finite-time adiabatic processes [19, 20], and by the technique of shortcut-to-adiabatically (STA) [21, 22, 23]. These two strategies lead to higher efficiencies and non-zero power.
In this paper, we have proposed a quantum harmonic heat engine constructed in an optomechanical system with an external mechanical parametric drive [24]. Unlike the other studies for QHEs in the optomechanical systems [25, 26, 27, 28, 29, 30, 31, 32, 33], the cavity is chosen as working substance. In particular, the mechanical mode has been considered as the quantum object that provides external energies, which is equivalent to the fuel as a component for classical machines. The heat flux for the total system in the cavity thermalization process is discussed, and the cyclic work extraction can be achieved from a single physical thermal reservoir coupled to the cavity. Due to the optomechanical interaction, non-equilibrium behavior is induced, resulting in this counter-intuitive phenomenon [34]. We then construct the Otto-type cycle for the heat engine by modulating the parametric drive strength [35]. Finally, the extra heat absorption induced by the optomechanical interaction in the work generation strokes has been considered, and the corresponding thermal efficiencies with and without the STA method are analyzed.
@@
We consider an optomechanical system with a parametric drive applied to mechanics [24, 35] and a coherent laser drive $H_{\text{dri}}=Aae^{i\omega_{L}t}+\text{h.c.}$, under the standard linearization, the system has the effective Hamiltonian
@@
where $H_{a}=\Delta(a^{\dagger}a+1/2)$, $H_{b}=\omega_{M}(b^{\dagger}b+1/2)$, $\Delta$ is the cavity detuning, $\omega_{M}$ is the intrinsic frequency for mechanical mode, $G$ is the optomechanical coupling strength, $\lambda$ indicates the parametric drive strength. The cavity mode is chosen as the quantum harmonic heat engine, with the detuning change to produce work. Here the cavity detuning $\Delta=\omega_{c}-\omega_{L}$ can be tuned in situ. In the weak interacting $G\ll\omega_{M}$ and resolved sideband regimes $\kappa\ll\omega_{M}$ [36, 37], the optomechanical interaction would only be resonantly enhanced, and the dynamics of the total optomechanical system can be described by the following master equation:
@@
$\begin{split}\mathcal{L}_{\text{diss}}\rho&=\kappa(\bar{n}_{c}+1)D[a]\rho+\kappa\bar{n}_{c}D[a^{\dagger}]\rho\\ &+\gamma(\bar{n}+1)D[b]\rho+\gamma\bar{n}D[b^{\dagger}]\rho,\end{split}$ (3)
where $D[o]\rho=o\rho o^{\dagger}-(o^{\dagger}o\rho+\rho o^{\dagger}o)/2$. Based on the second law of thermodynamics, a heat engine must be in contact with at least two thermal baths, with different thermal occupations $\bar{n}_{c}$ and decay rate $\kappa$ in corresponding thermodynamic processes. However, when an auxiliary term that interacts with the heat engine is taken into account, this interaction introduces an additional excitation to the heat engine, effectively acting as a hot (or cold) reservoir that couples solely with the heat engine. In our work, the optomechanical interaction and the mechanical parametric drive would excite the cavity into higher occupations at the non-equilibrium steady-states, and thus produce an effective hot thermal bath for the heat engine. Therefore, the mechanical mode with parametric drive here fuels the cavity, and the interaction term transports the energy from the fuel to the engine.
II.1 Heat flux with the parametric drive
As mentioned above, the mechanical mode acts as quantum fuel, providing effective hot reservoirs. Meanwhile, the cavity extracts heat from the mechanical mode via the optomechanical interaction. Naturally, the utilization efficiency of the energy of the quantum fuel, which can be defined as the ratio of heat flux transforming into the heat engine to the total heat flux injected into the mechanical mode, is one of the characteristics of quantum heat engines.
Based on energy conservation, the heat flux for both the cavity and the mechanical mode has the form [38]
@@ II.1 Heat flux with the parametric drive
The detailed derivation is given in Appendix A, here $\dot{Q}_{x}$ indicates the different contribution for the heat flux. As illustrated in Fig. 1(a), $\dot{Q}_{\rm{PD}}$ is the energy injection via mechanical parametric drive, $\dot{Q}_{b\rightarrow a}$ indicates heat flux transferred from the fuel to the heat engine, $\dot{Q}^{\uparrow}_{i}(i=a,b)$ is the heating flux from the reservoir, $\dot{Q}^{\downarrow}_{i}{(i=a,b)}$ is the dissipating flux. The exact form of these heat flux terms can be derived from the master equation(Eq. (2)),
@@
The definition of the utilization efficiency of energy for the mechanical mode is
@@
and the steady state occupation of photon mode is defined as $\bar{n}_{h}=\langle a^{\dagger}a\rangle_{ss}$ (Eq. (32)). The steady-state utilization efficiency of energy produced from the quantum fuel, and the effective thermal occupations for the cavity, are plotted in Fig. 1(b). Compared with the standard optomechanical interaction case ($\lambda=0$), the external mechanical driven term generates additional energy injection, which supplies more energy than just extracted from the mechanical thermal bath. From the view of quantum thermodynamics, the larger steady-state occupation for the cavity indicates higher inner energy differences, which reflects more heat absorption or work production. On the other hand, following the discussion of heat flux for the total system, the larger parametric drive strength leads to a higher utilization efficiency of energy, thus more heat is transmitted from the mechanical mode into the cavity. Therefore, the cavity has a stronger ability to extract heat from the additional effective heat bath. Both of the above qualities indicate that in this optomechanical model, the cavity is a suitable candidate for heat engines used to produce mechanical output work, while the mechanical mode, which injects the additional heat into the engine and creates an effective hot thermal reservoir, has been acting as the quantum fuel.
@@ II.2 Quantum Otto-type Cycle
A typical Otto cycle is constructed by two isochoric and two adiabatic processes. In the isochoric process, the engine is in contact with the hot(cold) thermal bath and has an energy exchange, while in the adiabatic process, the engine does not have thermal contact with the environment and just produces work. However, in quantum scenarios, the adiabatic processes not only satisfy the thermal adiabaticity that the system has zero thermal exchange but also follow the quantum adiabatic dynamics [3, 7, 5]. Therefore, we refer to the former case as the quantum Otto-type cycle.
@@
Figure 2: (a) The quantum Otto-type cycle for the cavity by tuning the detuning to produce the extra work output, and turning the parametric drive on and off to "heating" and "cooling" the cavity. (b) An illustration of the amplitude and detuning for the laser drive in an Otto-type cycle. The solid lines indicate detuning while the dashed lines are the corresponding amplitude. The black dashed-dotted line is $|\alpha|^{2}$, where $\alpha=A/(i\kappa-\Delta)$ is the classical part of the cavity mode. Thus the optomechanical coupling strength $G$ is fixed during the cycle because of $G\propto|\alpha|^{2}$.
(1). Isentropic compression($\rho_{1}\rightarrow\rho_{2}$), the cavity detuning $\Delta(t)$ and the amplitude of laser drive $A$ will be changed from ($\Delta_{l}$, $A_{l}$) to ($\Delta_{h}$, $A_{h}$) simultaneously during time $\tau$ to fix the classical part of the cavity mode. Thus in this stroke, the optomechanical coupling $G$ is fixed. During this stroke, $\lambda=0$.
@@
To design the two isentropic strokes, the evolution time $\tau\ll 1/\kappa$ is chosen to neglect the energy dissipation. Hence, in the large detuned case, $|\Delta_{l}-\Delta_{h}|\gg\kappa$, the steady state for the cavity in the cold isochoric stroke would almost be the thermal state with average occupation $\bar{n}_{c}$, and the optomechanical interaction term can be regarded as perturbation in analyzing the two isentropic strokes. Thus, the thermal efficiency in the absence of optomechanical interaction during the two isentropic processes can serve as a benchmark. Meanwhile, while the mechanical parametric amplification during the hot isochoric stroke does indeed generate squeezing for the cavity field, the steady state remains a Gaussian state (see Appendix B for the exact form of the steady-state P-function). Consequently, the thermal decomposition of a Gaussian state [39] indicates that, after the hot isochoric stroke, the eigenstate of the cavity mode follows a thermal distribution with an average occupation number $\bar{n}_{h}$.
The work production and heat absorption in a quantum Otto cycle can be defined with the two-point measurement approach, in which the energy change for the heat engine is obtained by projective measurement at the beginning and the end of each stroke [40]. The joint probability distribution for an Otto cycle has the form [17]
@@
Here $P^{\tau}_{n\rightarrow m}$ and $P^{\tau}_{k\rightarrow j}$ indicate the transition probabilities from initial eigenstates $n(k)$ to the final eigenstates $m(j)$ for the isentropic strokes [41, 12, 13], and $P_{n}(N)=(1/(1+N))(N/(1+N))^{n}$ is the probability distribution for the thermal states with occupation $N$.
For the calculation of average net work output and heat absorbed without optomechanical interaction in the isentropic strokes per cycle, the characteristic function of the joint probability distribution is
$\chi(u,v)=\int dWdQe^{iuW}e^{ivQ}P(W,Q)=\chi_{h}(u,v)\chi_{l}(u,v),$ (9)
@@
$\langle W\rangle_{Q^{*}}$ $=-i\frac{\partial\text{ln}\chi(u,v)}{\partial u}\mid_{u=v=0},$ (12a)
$\langle Q\rangle_{Q^{*}}$ $=-i\frac{\partial\text{ln}\chi(u,v)}{\partial v}\mid_{u=v=0}.$ (12b)
@@
$\eta_{\text{th}}=1-\frac{\Delta_{l}}{\Delta_{h}}\frac{Q^{*}(\bar{n}_{h}+1/2)-(\bar{n}_{c}+1/2)}{(\bar{n}_{h}+1/2)-Q^{*}(\bar{n}_{c}+1/2)}.$ (13)
@@ III Work generation and efficiencies
The quantum Otto engine reaches its maximal thermal efficiency when the isentropic strokes evolve quantum adiabaticity($Q^{*}=1$), with which the adiabatic Otto efficiency $\eta_{\rm{Qtto}}=1-\Delta_{l}/\Delta_{h}$. However, the quantum adiabatic condition requires the dynamical evolution of the system to be slow enough that the non-adiabatic transitions between different eigenstates can be neglected, and the system follows the instantaneous eigenstates in time. Therefore, even if the quantum adiabatic processes would have maximal efficiency, the power for the total work generation is zero. To obtain the efficiency close to the adiabatic case in finite time, the STA technique [23, 21] has been considered.
@@ III.1 Extra heat produced by optomechanical interaction
Figure 4: Efficiencies for different controlling parameters under fixed $\tau$. (a) Efficiencies for different detunings with a fixed ratio of $\Delta_{l}/\Delta_{h}=1/2$, where the mechanical parametric driving strength $\lambda=0.4\kappa$; (b) Efficiencies vs $\lambda$ with the same ratio of $\Delta_{l}/\Delta_{h}=1/2$ and fixed $\Delta_{h}=20\kappa$. In both plots, the evolution time has been chosen as $\kappa\tau=0.1$. Other parameters for both plots are chosen as $\bar{n}_{c}=0.01$, $\bar{n}=100$, $G=\kappa$, $\gamma=0.01\kappa$ and $\omega_{M}=\Delta_{h}$.
The thermal efficiency, when optomechanical coupling is taken into account, is given by $\eta^{\prime}_{\text{th}}=-\langle W\rangle_{Q^{*}}/(\langle Q\rangle_{Q^{*}}+\sum_{j=c,e}\langle\Delta E^{j}\rangle_{\tau})$. Here $Q^{*}>1$ because of the nonadiabatic nature, and the additional heat absorption $\langle\Delta E^{\rm{c/e}}\rangle_{\tau}$, resulting from the optomechanical interaction, has been included. The efficiencies for the fixed evolution time $\tau$ are plotted in Fig. 4. In both plots, the nonadiabatic efficiencies are calculated by Eq. (13), the modification for the heat absorption induced by the optomechanical interaction is added to the total heat absorption, and the detuning was tuned following the Eq. (16). The efficiencies for the STA methods that neglect and include the extra heat terms are calculated by Eq.(18) and Eq. (22), respectively. It has been noted that as $\Delta_{h}$ and $\Delta_{l}$ increase, the STA Hamiltonian(Eq. (14)) has less contribution to the system evolution, the system evolves near the instantaneous eigenstates once the boundary conditions are fulfilled. Meanwhile, because the contribution to the total heat absorption in a thermodynamic cycle has three parts, heat absorbing from the effective hot thermal reservoir $\langle Q\rangle$, the counter-diabatic driving $\langle H^{c/e}_{\rm{STA}}\rangle_{\tau}$ and the extra heat from the optomechanical interaction(Eq. (21)), the contribution for the latter two terms is mainly determined by the evolution time for the two strokes, thus the influence of these two parts decreases as mechanical parametric driving strength increases.